difference of two numbers is 24, larger number is four less than three times the smaller number
Answer by JulietG(1812)

A - B = 24 [difference of two numbers is 24]
A = 3B - 4 [larger number is four less than three times the smaller number]
.
Replace the known value of A from the second equation into the first.
(3B-4) - B = 24
Subtract
2B-4 = 24
Add 4 to each side
2B = 28
Divide each side by 2
B = 14
.
If the smaller number is 14 and the difference is 24, then A = 14+24, or 38
.
Let's see if that works:
38 = (3*14) - 4
38 = 42-4
38=38
Success!